# Multivariable Painlevé-II equation: connection formulas for asymptotic solutions

### 23 Mar 2026 v1

@@ II Definition of the model
Let a system of $n$ coupled nonlinear differential equations depend on $n$ parameters $e_{1}<e_{2}<\ldots<e_{n}$:
$u_{1}^{\prime\prime}(x)$ $=$ $xu_{1}(x)-2u_{1}(x)\sum_{k=1}^{n}u_{n}^{2}(x)+e_{1}u_{1}(x),$
$u_{2}^{\prime\prime}(x)$ $=$ $xu_{2}(x)-2u_{2}(x)\sum_{k=1}^{n}u_{n}^{2}(x)+e_{2}u_{2}(x),$ (2)
@@
$u_{n}^{\prime\prime}(x)$ $=$ $xu_{n}(x)-2u_{n}(x)\sum_{k=1}^{n}u_{n}^{2}(x)+e_{n}u_{n}(x).$
A shift, $x\rightarrow x-e_{1}$, sets $e_{1}=0$, which will be assumed here.
@@
$\left[\begin{array}[]{ccccc}4t^{2}+x-2\sum\limits_{k=1}^{n}u_{k}^{2}(x)&0&\cdots&\cdots&0\\ 0&-(4t^{2}+x+2e_{1})+2u_{1}^{2}&2u_{1}u_{2}&\cdots&2u_{1}u_{n}\\ \vdots&2u_{2}u_{1}&-(4t^{2}+x+2e_{2})+2u_{2}^{2}&\ddots&2u_{2}u_{n}\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&2u_{n}u_{1}(x)&2u_{n}u_{2}&\cdots&-(4t^{2}+x+2e_{n})+2u_{n}^{2}\end{array}\right],$ (15)
@@ III The system of two variables: main results
$u_{2}^{\prime\prime}(x)$ $\!=\!$ $xu_{2}(x)\!-\!2u_{2}(x)\left[u_{1}^{2}(x)\!+\!u_{2}^{2}(x)\right]\!-\!eu_{2}(x),$ (26)
where $e>0$.
@@ III.1 Connection formulas
$u_{2}$ $\!=\!$ $\frac{\alpha_{2}}{(-x+e)^{1/4}}\sin[\frac{2}{3}(-x+e)^{3/2}+$ (27)
@@
$u_{2}(x)$ $=$ $\sigma A\cos\left[\sqrt{e}x-\frac{A^{2}\sqrt{e}}{2}\ln x+\phi_{2}\right].$ (29)
@@
$I_{1}\equiv\frac{\rho^{2}}{2},\quad I_{2}\equiv\frac{A^{2}\sqrt{e}}{2},$ (30)
@@
$\Phi_{1}$ $\equiv$ $\frac{\pi}{4}+{\rm arg}\Gamma\left(i\frac{\alpha_{1}^{2}}{2}\right)+\varphi_{1}-\frac{3\alpha_{1}^{2}}{2}\ln 2+\frac{\alpha_{2}^{2}}{2}\ln(e/4),$
$\Phi_{2}$ $\equiv$ $\frac{\pi}{4}+{\rm arg}\Gamma\left(i\frac{\alpha_{2}^{2}}{2}\right)+\varphi_{2}-\frac{3\alpha_{2}^{2}}{2}\ln 2+\frac{\alpha_{1}^{2}}{2}\ln(e/4).$
The main result of this article is the following connection formulas that relate the final parameters $\{\sigma,I_{1},\,I_{2},\phi_{1},\phi_{2}\}$ to the set of initial parameters $\{p_{1},p_{2},\Phi_{1},\Phi_{2}\}$ and the parameter of the equation, $e$:
@@
$\phi_{2}$ $=$ $\frac{3\pi}{4}-\frac{2}{3}e^{3/2}-I_{2}\ln(4e^{1/2})+{\rm arg}\left[\Gamma(iI_{2})\right]-{\rm arg}\left(e^{i\Phi_{2}}+e^{-i\Phi_{2}}\left(p_{1}+(1-p_{1})e^{2i\Phi_{1}}\right)\right).$ (36)
Figure 1: Numerical test of Eqs. (33)-(36). Solid curves correspond to theoretical predictions for the dependence on the equation parameter $e\in(0.05,5)$ (a) of the action variables $I_{1}$ (brown) and $I_{2}$ (blue); (b) of the angle-related variables $\sin\phi_{1}$ (brown) and $\sin\phi_{2}$ (blue). Discrete points correspond to results of numerical simulations. Both for analytical predictions and numerical calculations the initial conditions were chosen to be $\alpha_{1}=0.9$, $\alpha_{2}=0.8$, $\varphi_{1}=\pi/2$, $\varphi_{2}=\pi/3$. Simulations were performed in the interval $x\in(-5000,5000)$ with a discretization step $dx=0.00001$. The algorithm is described in [31]. Both theory and numerical simulations agree on that $\sigma=-1$ for all points here.
Figure 2: Numerical test of Eqs. (32)-(36). Solid curves correspond to theoretical predictions for the dependence on the angle $\varphi_{1}\in(0,\pi)$ (a,b) of the action variables $I_{1}$ (brown) and $I_{2}$ (blue), respectively; (c,d) of the angle-related variables $\sin\phi_{1}$ (brown) and $\sin\phi_{2}$ (blue), respectively. Discrete points correspond to results of numerical simulations. Both for analytical predictions and numerical calculations the initial conditions were chosen to be $e=1$, $\alpha_{1}=0.8$, $\alpha_{2}=0.6$, $\varphi_{2}=\pi/2$. Simulations were performed in the interval $x\in(-5000,5000)$ with a discretization step $dx=0.00001$. Inset in (b) shows the test of Eq. (32) for the sign $\sigma$ dependence on $\varphi_{1}$ (purple).
@@ III.2 Tests of connection formulas
Finally, let me comment on corrections to the connection formulas. For large but finite $x$, they appear starting with the terms $\propto 1/\sqrt{x}$, which decay relatively slowly. They can be calculated using the same WKB approach, but the complete study of such terms appears to be difficult due to the large number of diverse relevant contributions. Some of them are singular in the limit $e\rightarrow 0$. This is why the simulations were restricted to $e>0.05$ – otherwise, a considerably larger integration interval is needed in order to achieve a comparable agreement with theoretical predictions.
Numerically, for the range of the tested parameters, with $e>0.05$, there were two specific corrections observed that dominated over other terms of the same order in $x$. These two corrections can be easily included into the connection formulas, leading to noticeable accuracy improvements at large but finite $x$. First, the regular part of $u_{1}(x)$ in Eq. (III.1) is renormalized by interactions as
@@
$\phi_{1}\rightarrow\phi_{1}-2\pi I_{2}\sqrt{\frac{e}{2x}}.$ (38)
Although well suppressed, small effects of this correction were visible even for the longest simulations. This is attributed to its growth with $e$ and the existence of singularities that make $I_{2}$ anomalously large at specific initial conditions.
@@ IV.2 WKB analysis in the limit $x\rightarrow$-$\infty$
$\small{H(\tau)=\left[\begin{array}[]{ccc}|x|^{3/2}(4\tau^{2}-1)-2\sqrt{x}(u_{1}^{2}(x)+u_{2}^{2}(x))&-4|x|\tau u_{1}-2i\sqrt{|x|}u_{1}^{\prime}&-4|x|\tau u_{2}-2i\sqrt{|x|}u_{2}^{\prime}\\ -4|x|\tau u_{1}+2i\sqrt{|x|}u_{1}^{\prime}&-|x|^{3/2}(4\tau^{2}-1)+2\sqrt{|x|}u_{1}(x)^{2}&2\sqrt{|x|}u_{1}(x)u_{2}(x)\\ -4|x|\tau u_{2}+2i\sqrt{|x|}u_{2}^{\prime}&2\sqrt{|x|}u_{1}(x)u_{2}(x)&-|x|^{3/2}(4\tau^{2}-1)+2\sqrt{|x|}(e+u_{2}(x)^{2})\end{array}\right].}$ (45)
Figure 4: (a) Eigenvalues of the Hamiltonian (45) as functions of $t=\sqrt{|x|}\tau$ at large negative $x$ (blue curves). Here, x=16x=-16; $\alpha_{1}=\alpha_{2}=0.1$, $\varphi_{1}=\pi/3$, $\varphi_{2}=3\pi/4$. The values of $u_{1}(x)$ and $u_{2}(x)$ are approximated by asymptotic formulas in Eq. (III.1). The Hamiltonian (45) is written in the basis of diabatic states. The labels 0, 1, 20,\,1,\,2 define the convention for the diabatic state indices. The diabatic states coincide asymptotically with the eigenstates as $t\rightarrow\pm\infty$. The red and green arrows show two semiclassical trajectories that originate on diabatic level 0 as $t\rightarrow$-$\infty$ and terminate at level 1 as $t\rightarrow+\infty$. The dashed circle encloses the behavior of the energy levels near the pseudo-time point $t=\sqrt{|x|}/2$. (b) Near $t=\sqrt{|x|}/2$, the dynamics is described by the exactly solvable DOM for three levels that cross linearly at fixed couplings $g_{1}^{+}$ and $g_{2}^{+}$. The diagram of diabatic levels shows only the time dependence of diagonal elements of the Hamiltonian, which in the DOM are straight lines in the time-energy plot. (c) Eigenvalues of the Hamiltonian $H(t,x)+x^{3/2}(4\tau^{2}+1)\hat{1}$, as functions of $t=\sqrt{x}\tau$ at large positive $x$ (blue). Here, $x=10$, $A=0.2$, $\rho=0.12$, $e=1.2$. A term proportional to $\hat{1}$ was added to the Hamiltonian to expose the spectrum better. The values of $u_{1}(x)$ and $u_{2}(x)$ are approximated by asymptotic formulas in Eqs. (III.1) and (29), respectively. The red arrows show the unique path that connects diabatic levels 0 and $1$. At large positive $x$, it passes through two avoided crossing points near times $t_{0}$ and $t_{+}$.
@@
$H_{-1/2}(\tau_{m})=\left[\begin{array}[]{ccc}-4|x|^{3/2}\tau_{m}&g_{1}^{-}&g_{2}^{-}\\ (g_{1}^{-})^{*}&4|x|^{3/2}\tau_{m}&0\\ (g_{2}^{-})^{*}&0&4|x|^{3/2}\tau_{m}+2\sqrt{|x|}e\end{array}\right].$ (46)
@@
$H_{1/2}(\tau_{p})=\left[\begin{array}[]{ccc}4|x|^{3/2}\tau_{p}&g_{1}^{+}&g_{2}^{+}\\ (g_{1}^{+})^{*}&-4|x|^{3/2}\tau_{p}&0\\ (g_{2}^{+})^{*}&0&-4|x|^{3/2}\tau_{p}+2\sqrt{|x|}e\end{array}\right].$ (47)
@@
This WKB approach is equivalent to the textbook independent crossing approximation, which is widely used in theoretical physics and described, e.g., in Section 5.4 of [29]. This version of WKB was used in the derivation of transition amplitudes in multistate Landau-Zener models [7]. The advantage of this approach over the WKB analysis commonly used in studies of Painlevé equations [12] is that the pseudo-time $t$ is never treated as a complex variable, so there is no need to analyze the full Stokes phenomenon to obtain the evolution matrix $U_{P}$.
In the present case, the full matrix $U_{P}$ is not required. It is sufficient to assume that initially the system is in state $|0\rangle$, which has the highest energy as $t\rightarrow$-$\infty$, thus leading to the derivation of transition amplitudes $U_{P,00}$, $U_{P,10}$, and $U_{P,20}$.
@@ IV.3 WKB analysis as $x\rightarrow+\infty$
This limit is simpler because, as illustrated in Fig. 4(c), the level crossing pattern leads only to three elementary avoided crossings between pairs of levels. These crossings are well separated both in energy and in time $t$. For example, the times of diabatic level crossings are given by $t_{0}=0$ and $t_{\pm}\approx\pm x/(4\sqrt{e})$. Therefore, each crossing is described by the scattering matrix for the two-state Landau-Zener model. Moreover, for the initial conditions with only level 0 populated, the avoided crossing at $t_{-}$ is irrelevant, so only a single trajectory connects the initial state $|0\rangle$ to any final diabatic state. Therefore, path interference does not occur. Figure 4(c) illustrates this property by showing the only semiclassical trajectory connecting levels 0 and $1$.

### 27 Mar 2026 v2

@@ II Definition of the model
Let a system of $n$ coupled nonlinear differential equations depend on $n$ parameters $\varepsilon_{1}<\varepsilon_{2}<\ldots<\varepsilon_{n}$:
$u_{1}^{\prime\prime}(x)$ $=$ $xu_{1}(x)-2u_{1}(x)\sum_{k=1}^{n}u_{n}^{2}(x)-\varepsilon_{1}u_{1}(x),$
$u_{2}^{\prime\prime}(x)$ $=$ $xu_{2}(x)-2u_{2}(x)\sum_{k=1}^{n}u_{n}^{2}(x)-\varepsilon_{2}u_{2}(x),$ (2)
@@
$u_{n}^{\prime\prime}(x)$ $=$ $xu_{n}(x)-2u_{n}(x)\sum_{k=1}^{n}u_{n}^{2}(x)-\varepsilon_{n}u_{n}(x).$
A shift, $x\rightarrow x+\varepsilon_{1}$, sets $\varepsilon_{1}=0$, which will be assumed here.
@@
$\left[\begin{array}[]{ccccc}4t^{2}+x-2\sum\limits_{k=1}^{n}u_{k}^{2}(x)&0&\cdots&\cdots&0\\ 0&-(4t^{2}+x-2\varepsilon_{1})+2u_{1}^{2}&2u_{1}u_{2}&\cdots&2u_{1}u_{n}\\ \vdots&2u_{2}u_{1}&-(4t^{2}-2\varepsilon_{2})+2u_{2}^{2}&\ddots&2u_{2}u_{n}\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&2u_{n}u_{1}(x)&2u_{n}u_{2}&\cdots&-(4t^{2}+x-2\varepsilon_{n})+2u_{n}^{2}\end{array}\right],$ (15)
@@ III The system of two variables: main results
$u_{2}^{\prime\prime}(x)$ $\!=\!$ $xu_{2}(x)\!-\!2u_{2}(x)\left[u_{1}^{2}(x)\!+\!u_{2}^{2}(x)\right]\!-\!\varepsilon u_{2}(x),$ (26)
where $\varepsilon>0$.
@@ III.1 Connection formulas
$u_{2}$ $\!=\!$ $\frac{\alpha_{2}}{(-x+\varepsilon)^{1/4}}\sin[\frac{2}{3}(-x+\varepsilon)^{3/2}+$ (27)
@@
$u_{2}(x)$ $=$ $\sigma A\cos\left[\sqrt{\varepsilon}x-\frac{A^{2}\sqrt{\varepsilon}}{2}\ln x+\phi_{2}\right].$ (29)
@@
$I_{1}\equiv\frac{\rho^{2}}{2},\quad I_{2}\equiv\frac{A^{2}\sqrt{\varepsilon}}{2},$ (30)
@@
$\Phi_{1}$ $\equiv$ $\frac{\pi}{4}+{\rm arg}\Gamma\left(i\frac{\alpha_{1}^{2}}{2}\right)+\varphi_{1}-\frac{3\alpha_{1}^{2}}{2}\ln 2+\frac{\alpha_{2}^{2}}{2}\ln(\varepsilon/4),$
$\Phi_{2}$ $\equiv$ $\frac{\pi}{4}+{\rm arg}\Gamma\left(i\frac{\alpha_{2}^{2}}{2}\right)+\varphi_{2}-\frac{3\alpha_{2}^{2}}{2}\ln 2+\frac{\alpha_{1}^{2}}{2}\ln(\varepsilon/4).$
The main result of this article is the following connection formulas that relate the final parameters $\{\sigma,I_{1},\,I_{2},\phi_{1},\phi_{2}\}$ to the set of initial parameters $\{p_{1},p_{2},\Phi_{1},\Phi_{2}\}$ and the parameter of the equation, $\varepsilon$:
@@
$\phi_{2}$ $=$ $\frac{3\pi}{4}-\frac{2}{3}\varepsilon^{3/2}-I_{2}\ln(4\varepsilon^{1/2})+{\rm arg}\left[\Gamma(iI_{2})\right]-{\rm arg}\left(e^{i\Phi_{2}}+e^{-i\Phi_{2}}\left(p_{1}+(1-p_{1})e^{2i\Phi_{1}}\right)\right).$ (36)
Figure 1: Numerical test of Eqs. (33)-(36). Solid curves correspond to theoretical predictions for the dependence on the equation parameter $\varepsilon\in(0.05,5)$ (a) of the action variables $I_{1}$ (brown) and $I_{2}$ (blue); (b) of the angle-related variables $\sin\phi_{1}$ (brown) and $\sin\phi_{2}$ (blue). Discrete points correspond to results of numerical simulations. Both for analytical predictions and numerical calculations the initial conditions were chosen to be $\alpha_{1}=0.9$, $\alpha_{2}=0.8$, $\varphi_{1}=\pi/2$, $\varphi_{2}=\pi/3$. Simulations were performed in the interval $x\in(-5000,5000)$ with a discretization step $dx=0.00001$. The algorithm is described in [32]. Both theory and numerical simulations agree on that $\sigma=-1$ for all points here.
Figure 2: Numerical test of Eqs. (32)-(36). Solid curves correspond to theoretical predictions for the dependence on the angle $\varphi_{1}\in(0,\pi)$ (a,b) of the action variables $I_{1}$ (brown) and $I_{2}$ (blue), respectively; (c,d) of the angle-related variables $\sin\phi_{1}$ (brown) and $\sin\phi_{2}$ (blue), respectively. Discrete points correspond to results of numerical simulations. Both for analytical predictions and numerical calculations the initial conditions were chosen to be $\varepsilon=1$, $\alpha_{1}=0.8$, $\alpha_{2}=0.6$, $\varphi_{2}=\pi/2$. Simulations were performed in the interval $x\in(-5000,5000)$ with a discretization step $dx=0.00001$. Inset in (b) shows the test of Eq. (32) for the sign $\sigma$ dependence on $\varphi_{1}$ (purple).
@@ III.2 Tests of connection formulas
Finally, let me comment on corrections to the connection formulas. For large but finite $x$, they appear starting with the terms $\propto 1/\sqrt{x}$, which decay relatively slowly. They can be calculated using the same WKB approach, but the complete study of such terms appears to be difficult due to the large number of diverse relevant contributions. Some of them are singular in the limit $\varepsilon\rightarrow 0$. This is why the simulations were restricted to $\varepsilon\geq 0.05$ – otherwise, a considerably larger integration interval is needed in order to achieve a comparable agreement with theoretical predictions.
Numerically, for the range of the tested parameters, with $\varepsilon\geq 0.05$, there were two specific corrections observed that dominated over other terms of the same order in $x$. These two corrections can be easily included into the connection formulas, leading to noticeable accuracy improvements at large but finite $x$. First, the regular part of $u_{1}(x)$ in Eq. (III.1) is renormalized by interactions as
@@
$\phi_{1}\rightarrow\phi_{1}-2\pi I_{2}\sqrt{\frac{\varepsilon}{2x}}.$ (38)
Although well suppressed, small effects of this correction were visible even for the longest simulations. This is attributed to its growth with $\varepsilon$ and the existence of singularities that make $I_{2}$ anomalously large at specific initial conditions.
@@ IV.2 WKB analysis in the limit $x\rightarrow$-$\infty$
$\small{H(\tau)=\left[\begin{array}[]{ccc}|x|^{3/2}(4\tau^{2}-1)-2\sqrt{x}(u_{1}^{2}(x)+u_{2}^{2}(x))&-4|x|\tau u_{1}-2i\sqrt{|x|}u_{1}^{\prime}&-4|x|\tau u_{2}-2i\sqrt{|x|}u_{2}^{\prime}\\ -4|x|\tau u_{1}+2i\sqrt{|x|}u_{1}^{\prime}&-|x|^{3/2}(4\tau^{2}-1)+2\sqrt{|x|}u_{1}(x)^{2}&2\sqrt{|x|}u_{1}(x)u_{2}(x)\\ -4|x|\tau u_{2}+2i\sqrt{|x|}u_{2}^{\prime}&2\sqrt{|x|}u_{1}(x)u_{2}(x)&-|x|^{3/2}(4\tau^{2}-1)+2\sqrt{|x|}(\varepsilon+u_{2}(x)^{2})\end{array}\right].}$ (45)
Figure 4: (a) Eigenvalues of the Hamiltonian (45) as functions of $t=\sqrt{|x|}\tau$ at large negative $x$ (blue curves). Here, x=16x=-16; $\alpha_{1}=\alpha_{2}=0.1$, $\varphi_{1}=\pi/3$, $\varphi_{2}=3\pi/4$. The values of $u_{1}(x)$ and $u_{2}(x)$ are approximated by asymptotic formulas in Eq. (III.1). The Hamiltonian (45) is written in the basis of diabatic states. The labels 0, 1, 20,\,1,\,2 define the convention for the diabatic state indices. The diabatic states coincide asymptotically with the eigenstates as $t\rightarrow\pm\infty$. The red and green arrows show two semiclassical trajectories that originate on diabatic level 0 as $t\rightarrow$-$\infty$ and terminate at level 1 as $t\rightarrow+\infty$. The dashed circle encloses the behavior of the energy levels near the pseudo-time point $t=\sqrt{|x|}/2$. (b) Near $t=\sqrt{|x|}/2$, the dynamics is described by the exactly solvable DOM for three levels that cross linearly at fixed couplings $g_{1}^{+}$ and $g_{2}^{+}$. The diagram of diabatic levels shows only the time dependence of diagonal elements of the Hamiltonian, which in the DOM are straight lines in the time-energy plot. (c) Eigenvalues of the Hamiltonian $H(t,x)+x^{3/2}(4\tau^{2}+1)\hat{1}$, as functions of $t=\sqrt{x}\tau$ at large positive $x$ (blue). Here, $x=10$, $A=0.2$, $\rho=0.12$, $\varepsilon=1.2$. A term proportional to $\hat{1}$ was added to the Hamiltonian to expose the spectrum better without changing its essential features. The values of $u_{1}(x)$ and $u_{2}(x)$ are approximated by asymptotic formulas in Eqs. (III.1) and (29), respectively. The red arrows show the unique path that connects diabatic levels 0 and $1$. At large positive $x$, it passes through two avoided crossing points near times $t_{0}$ and $t_{+}$.
@@
$H_{-1/2}(\tau_{m})=\left[\begin{array}[]{ccc}-4|x|^{3/2}\tau_{m}&g_{1}^{-}&g_{2}^{-}\\ (g_{1}^{-})^{*}&4|x|^{3/2}\tau_{m}&0\\ (g_{2}^{-})^{*}&0&4|x|^{3/2}\tau_{m}+2\sqrt{|x|}\varepsilon\end{array}\right].$ (46)
@@
$H_{1/2}(\tau_{p})=\left[\begin{array}[]{ccc}4|x|^{3/2}\tau_{p}&g_{1}^{+}&g_{2}^{+}\\ (g_{1}^{+})^{*}&-4|x|^{3/2}\tau_{p}&0\\ (g_{2}^{+})^{*}&0&-4|x|^{3/2}\tau_{p}+2\sqrt{|x|}\varepsilon\end{array}\right].$ (47)
@@
In the present case, the full evolution matrix $U_{P}$ is not required. It is sufficient to assume that initially the system is in state $|0\rangle$, which has the highest energy as $t\rightarrow$-$\infty$, thus leading to the derivation of transition amplitudes $U_{P,00}$, $U_{P,10}$, and $U_{P,20}$.
This WKB approach has been used to derive some of the earliest connection formulas for Painlevé equations [22]. Here, due to multiple encountered avoided crossings of the Landau-Zener type, it becomes equivalent to the textbook independent crossing approximation, which is also used in theoretical physics as approximation and for derivation of transition amplitudes in multistate Landau-Zener models [7] (author recommends Section 5.4 of Ref. [30] for elementary textbook introduction). The convenience of this approach is that the pseudo-time $t$ is never treated as a complex variable to obtain the evolution matrix elements. However, a mathematically more rigorous approach is based on solving the Riemann-Hilbert problem [12], which was not used in the present work, so the numerical simulations rather than more detailed mathematical analysis validate the connection formulas yet.
@@ IV.3 WKB analysis as $x\rightarrow+\infty$
This limit is simpler because, as illustrated in Fig. 4(c), the level crossing pattern leads only to three elementary avoided crossings between pairs of levels. These crossings are well separated both in energy and in time $t$. For example, the times of diabatic level crossings are given by $t_{0}=0$ and $t_{\pm}\approx\pm x/(4\sqrt{\varepsilon})$. Therefore, each crossing is described by the scattering matrix for the two-state Landau-Zener model. Moreover, for the initial conditions with only level 0 populated, the avoided crossing at $t_{-}$ is irrelevant, so only a single trajectory connects the initial state $|0\rangle$ to any final diabatic state. Therefore, path interference does not occur. Figure 4(c) illustrates this property by showing the only semiclassical trajectory connecting levels 0 and $1$.